OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 A. (14, 12) B. There are an infinite number of solutions. C. (12, 14) D. There is no solution.

OpenStudy (anonymous):

i am thinking either c or b

OpenStudy (anonymous):

let me work it out so i am sure one sec, because i am leaning towards c

OpenStudy (anonymous):

i have one more question after this and that is it just to make it clear, and i do think it is c :)

OpenStudy (anonymous):

So to substitute you first must simplify / rearrange one of the equations. In this case it is already rearranged nicely. Then, substitute in your answer from above. e.g. 6x - 2 (3x - 7) = 12 6x - 6x - 7 = 12 -7 = 12 The above is obviously not true, and so no matter what values you give x and y, it will not work. Therefore, there is no solution. D.

OpenStudy (anonymous):

aww really i was way off!

OpenStudy (anonymous):

thx here is a medal!

OpenStudy (anonymous):

You can always check something like C by putting the values into the equation. e.g. (12, 14) y = 3x – 7 does 14 = 3(12) - 7? 14 = 36 - 7 14 = 29 No, so it cannot be a solution